For the standard normal distribution, determine the following probabilities:
(a) Pr(Z  ≥  1.5)
(b) Pr(1.2  ≤  Z  ≤  1.75)
Answer #1

Givent for the standard Normal distribution we need to find (a) A(Z > 15) - 1-P (721.5) = 1-0.9332 Normeel table) = 0.0668 1-